OpenStudy (anonymous):

Write the equation in standard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (zzr0ck3r):

1/2x + b = y 1/2(5) + b = 3 b = 3-5/2 b =1/2 y = 1/2x + 1/2